### v2

@@ Abstract
This paper explores the potential for leveraging Large Language Models (LLM) in the realm of online advertising systems. We delve into essential requirements including privacy, latency, reliability as well as the satisfaction of users and advertisers which such a system must fulfill. We further introduce a general framework for LLM advertisement, consisting of modification, bidding, prediction, and auction modules. Different design considerations for each module is presented, with an in-depth examination of their practicality and the technical challenges inherent to their implementation. Finally, we explore the prospect of LLM-based dynamic creative optimization as a means to significantly enhance the appeal of advertisements to users and discuss its additional challenges.
@@ 1 Introduction
In the vast landscape of online search engines, the role of advertisements has become pivotal, shaping the digital experience for users globally. The enormity of the market, with trillions of dollars at play, underscores the economic significance of advertising. In a world where people increasingly expect free access to information and services, advertisements play a crucial role in subsidizing these offerings. The democratization of education, information, and other resources is made possible through ad-driven models, enabling individuals to not only consume content for free but also contribute to the digital ecosystem. Additionally, the symbiotic relationship between advertisements and content creation creates a feedback loop, fostering economic growth. Several online platforms often offer a subscription model for people to disable the advertisement in using their services, however, the advertisement market is becoming increasingly dominant, e.g., Netflix recently started a ad-supported plan Netflix (2022).
@@
Recently, large language models (LLMs) have gained widespread adoption among users, serving various functions including question answering, content generation, translation, code completion, and more Brown et al. (2020); Nijkamp et al. (2022); Fried et al. (2022); Wang and Komatsuzaki (2021); Liu et al. (2023). The proliferation of AI-driven assistant language models, such as ChatGPT, has contributed to a growing trend wherein individuals increasingly use such models to address their inquiries, occasionally replacing traditional search engines as their primary information-seeking tool. The substantial usage volumes stemming from diverse users shave prompted companies offering these tools, which we call LLM providers, to contemplate revenue generation through advertising AdWeek (2023); Crunch (2023); Microsoft (2023). Consequently, an interesting and fundamental question arises:
How can LLM providers generate revenue through running an online advertisement?
@@
The concept of online advertising has been extensively studied within the realm of search engines, where auctions are conducted among advertisements from advertisers when a user inputs a query. We focus on the prospect of transposing this online advertising model and auction framework to the context of large language models, exploring the potential challenges in this context. In this paper, our primary focus revolves around discussing different criteria for LLM advertisement and evaluating different frameworks based on them.
@@
Figure 1: An example of providing unstructured advertisement in the LLM. Left column refers to the case where we ask above question from ChatGPT 3.5, and to incorporate the ads we use queries specified in the right column.
@@
We start with illustrative scenarios where a user asks a technical question about computers (Figure 1). Without advertisement, LLM would typically generate a response to address the user’s query. On the other hand, within the content of the generated response, there exists an opportunity for the LLM provider to strategically incorporate advertisements for computer manufacturers. There is a spectrum of possibilities for including ad content, ranging from ads alongside the output, explicitly putting advertisements in the output like display ads, and including advertisements within the generated output. Also, in case the advertisements are included within the generated output, there are several possibilities such that the ads may simply replace one of the element in case they are written as a list of elements (e.g., Figure 4, denoted by structured output), or the ads itself is incorporated within the original text/figure without hurting/changing the context of the original output (e.g., Figure 1, denoted by unstructured output).
We will focus on the case where the ads are incorporated in the unstructured output of the LLM, since the other cases could be handled using techniques from modern online search/display advertisement system. Moreover, this kind of advertisement can be more interesting for both users and advertisers due to the huge capabilities of language models in generating high-quality responses while promoting products and services from advertisers. [^2] Indeed, it’s worth noting that Microsoft Bing has been running such types of advertisements in its AI chatbot Microsoft (2023). Then, a natural question is how the system can incorporate advertisements into the generated output to satisfy both the advertisers and the user. Further, it is not trivial how the advertisers should communicate with the system to bid on each modified output.
@@
In this paper, we discuss several desiderata for operating an online advertisement over LLM and propose a general framework to meet the requirements. Our framework consists of several modules: (i) modification in which the original output of LLM is modified; (ii) bidding that advertisers utilize to bid on the modified outputs; (iii) prediction in which the LLM provider computes required information about advertisements; and (iv) auction in which the advertises compete and the final output is selected. We discuss different designs and technical challenges of each of those modules, along with other technical challenges to implement the system in a practical manner.  [^3] We first focus on presenting a single advertisement in the LLM output, however, generalization of the proposed framework to incorporate multiple ads at once is discussed in Appendix B.2. Finally, in Section 6, as a LLM’s (more generally generative model’s) usecase of independent interest, we discuss how LLM can empower the online advertisement system to tailor the content of ads itself to individual users, leveraging their specific contexts to generate more appealing content, which is called as dynamic creative optimization (DCO) or responsive ad in the literature.
@@ 2 Related works
Here we present related works in terms of LLM, online advertisement system, and their intersection. Further related works can be found in Appendix A.
@@
Advancements in AI, NLP, and conversational agents, driven by Transformer architecture (Vaswani et al., 2017), have given rise to models like GPT-3 (Brown et al., 2020) and BERT (Devlin et al., 2019). These models revolutionize chatbots, enabling context-aware, human-like interactions across diverse domains Abd-Alrazaq et al. (2020); Nicolescu and Tudorache (2022).
@@ 3 Requirements
In this section, we outline our expectations for an LLM advertisement system, which should be evaluated based on the following criteria:
@@
LLM advertisement systems need to associate with advertisers, making the entire system vulnerable to harmful behaviors from them. To mitigate that, LLM providers should consider all aspects of the behavior of the advertisers, ensuring the reliability and alignment of the system (Hendrycks et al., 2020).
@@
When an advertisement is added to LLM output, it’s crucial to maintain a high level of quality in the generated content. Users do not wish to encounter an excessive amount of advertising in the generated responses, especially when it doesn’t align with their search or interests. Advertisement involvement should not significantly degrade LLM output, as doing so would lead to reduced user satisfaction, thereby decreasing the user retention. For instance, in modern online advertisement system, this is often handled using floor prices to filter out irrelevant ads to maintain the quality of the user experience.
@@
As advertisers pay LLM providers to involve advertisements in their outputs, they expect appropriate exposure of ads in them. Whether their products, services, or any other form of advertising, they expect these items to be showcased in a compelling and engaging manner within the output. This presentation should be interesting for users, ultimately driving revenue for the advertisers.
LLM provider’s revenue
@@
Obtaining revenue is the primary goal of providing LLM advertisement services, thus, LLM providers should ensure that the total additional cost of advertisement is compensated by the revenue obtained from advertisers. It is worth noting that advertisements may potentially reduce the overall number of users engaging with the system, which could have adverse effects on LLM providers.
4 Framework for LLM Advertisement
Figure 2: Overall framework of LAS.
@@ 4 Framework for LLM Advertisement
In this section, we present overall framework for LLM advertisement system (henceforth LAS). We introduce several feasible types of architectures for LAS, each of which has its own advantages and disadvantages that will be discussed together. Mainly, we focus on a scenario in which a single user provides a single query $q$ to the LLM, and the original output by LLM is given by $X$. Further, a context $c$ captures a variety of features that are relevant to the advertisement recommendation, e.g., history of the previous queries, user segment, region, and date. Although the number of advertisers (bidders) varies from time to time, at the moment the user inputs the query $q$, we suppose that there are $n$ advertisers indexed by $V=\{\text{adv}_{1},\text{adv}_{2},\ldots,\text{adv}_{n}\}$. For simplicity, we assume that each advertiser $\text{adv}_{i}$ is equipped with a single advertisement (ad) he/she wants to post. We denote $\text{adv}_{i}$’s ad by $\text{ad}_{i}$ for $i\in[n]$, and let $D=\{\text{ad}_{1},\ldots,\text{ad}_{n}\}$.
@@
Overall, we divide the LAS as follows based on their functionalities [^6] These functionalities may be distributed to multiple market players such as supply side platform, display side platform, or ad exchange, as does in the current online ad eco-system. into $4$ modules: (i) output modification, (ii) bidding, (iii) prediction, and (iv) auction. This overall framework is illustrated in Figure 2.
In short, the user asks a query $q$ to the system along with a context $c$, and the LAS first generates the original output $X$. Then, the modification module generates specific output $X_{i}$ for each of the advertisers, and the bidding module generates corresponding bid $b_{i}$. At the same time, the prediction module predicts the quality of the output $X_{i}$ by measuring the user satisfaction rate $\text{sr}_{i}$ and click-through-rate $\text{ctr}_{i}$, which will be elaborated later. Finally, the auction module runs auction to determine the final output and corresponding payment amount to charge the advertiser. Also, note that we do not restrict the final output to be exactly belonging to one of $X_{i}$, but can be totally different one based on the advertisers’ bids $b_{i}$ and corresponding output $X_{i}$. For the rest of the section, we will explain each module and its responsibility/functionality in a sequential manner.
Modification module
@@
Modification module takes responsibility for generating modified output based on corresponding ads. This module takes the pair of $(q,X,c)$ and a set of advertisements $D$ as the input, and returns $(X_{i})_{i\in[n]}$ where $X_{i}$ denotes the modified output for $\text{ad}_{i}$.
@@
Overall, we consider two approaches based on who generates the modified outputs. 777 For either cases, one can generate the modified output by throwing an additional query to the LLM as presented in Figure 1.
@@
    In the LAS modification model, LAS directly generates the modified output.
Bidding module
Bidding module generates bid based on the modified outputs. The input to the bidding module is the query $q$, context $c$, and modified outputs $(X_{i})_{i\in[n]}$, and the output will be a bid $(b_{i})_{i\in[n]}$, where $b_{i}$ is the bid for $ad_{i}$. Similar to the generation model, we consider two models for the bidding module.
@@
Since the static bidding model may operated solely based on some keywords that are relevant to the advertisement products, we often refer this to as keyword bidding. In Section 5, we discuss several extensions of these bidding models, thereby allowing more flexibility for advertisers to customize their advertisement strategy by bidding.
@@
Dynamic bidding model may be of interest to the advertisers who have their own measures to estimate the quality of the modified output and thus want to customize their bid in a way that might not be captured by the static bidding. For example, in extreme, suppose that LAS may only consider click-through-rate of the modified output and does not care about the user satisfaction rate or how much the inserted advertisement relates to the context of the output. In this case, the advertiser may want to customize their bid by measuring the quality of the output by itself, since he may be afraid of the system inserting its ads in an awkward manner.
Prediction Module

### v3

@@ Abstract
This paper explores the potential for leveraging Large Language Models (LLM) in the realm of online advertising systems. We delve into essential requirements including privacy, latency, reliability as well as the satisfaction of users and advertisers that such a system must fulfill. We further introduce a general framework for LLM advertisement, consisting of modification, bidding, prediction, and auction modules. Different design considerations for each module are presented. Fundamental questions regarding practicality, efficiency, and implementation challenges of these designs are raised for future research. Finally, we explore the prospect of LLM-based dynamic creative optimization as a means to significantly enhance the appeal of advertisements to users and discuss its additional challenges.
@@ 1 Introduction
In the vast landscape of online search engines, the role of advertisements has become pivotal, shaping the digital experience for users globally. The enormity of the market, with trillions of dollars at play, underscores the economic significance of advertising, e.g., the market size of search advertisement as of $2020$ was valued at USD $164.12$ billion (Insight, 2022).
@@
Online advertising not only brings revenue to platform companies, but also plays a crucial role in subsidizing the free access to information and service to every individual. The democratization of education, information, and other resources is made possible through ad-driven models, enabling individuals to not only consume content for free but also contribute to the digital ecosystem. Additionally, the symbiotic relationship between advertisements and content creation creates a feedback loop, fostering economic growth. As a side note, even the subscription-based streaming platforms like Netflix are starting ad-supported plans (Netflix, 2022).
@@
On the other hand, recently, large language models (LLMs) (Brown et al., 2020; Anil et al., 2023; Thoppilan et al., 2022) have gained widespread adoption among users, serving various functions including question answering, content generation, translation, code completion, and more (Nijkamp et al., 2022; Fried et al., 2022; Wang and Komatsuzaki, 2021; Liu et al., 2023a). The proliferation of AI-driven assistant language models, such as ChatGPT, has contributed to a growing trend wherein individuals increasingly use such models to address their inquiries, occasionally replacing traditional search engines as their primary information-seeking tool. According to PCMag (2023), even for now, $35\%$ of casual users say they find LLMs to be more helpful in finding information than search engine. It is obvious that such a trend will be accelerating in the near future as well. The substantial usage volumes stemming from diverse users would induce companies offering these tools, which we call LLM providers, to contemplate revenue generation through advertising (AdWeek, 2023; Crunch, 2023; Microsoft, 2023). Consequently, an interesting and fundamental question arises:
How can LLM providers make revenue by running an online advertisement?
@@
The concept of online advertising has been extensively studied within the realm of search engines, where auctions are conducted among advertisements from advertisers when a user inputs a query. This paper focuses on the prospect of transposing this online advertising model and auction framework to the context of large language models. We further discuss technical challenges and potential framework to run online advertisement system in LLM, thereby calling academic and industrial researchers to the area of importance.
@@
Figure 1: An example of providing unstructured advertisement in the LLM. Left column refers to the case where we ask above question from ChatGPT 4, and to incorporate the ads we use queries specified in the right column.
Search advertising
To better explain fundamental differences between the standard search advertising (SA) [^2] Its mechanism design problem is often called as sponsored search auction (SSA). and the LLM advertising (LLMA), we briefly introduce how the standard SA works (Lahaie et al., 2007) in what follows. (1) Bidding: in the SA, the owner of each ad $i$ writes bid $b_{i}\in\mathbb{R}_{\geq 0}$ on targeting keyword for $i\in[n]$, which can possibly a set of keywords. (2) Output generation: the platform first decides how many slots to allocate for ads in the search engine results page (SERP), say $k$. (3) Prediction: given $k$ slots in SERP, the platform then predicts the click-through-rate (CTR) $\alpha_{ij}$ when ad $i$ is allocated in slot $j$. (4) Auction: then it optimizes
$\sum_{i=1}^{n}\sum_{j=1}^{k}\alpha_{ij}b_{i}x_{ij},$ (1.1)
where $x_{ij}=1$ if ad $i$ is allocated in slot $j$, given the constraint $\sum_{i=1}^{n}x_{ij}\leq 1$ for every $j\in[k]$, and charges each ad with respect to pre-committed payment rule.
Overall, whenever a user arrives in the platform and searches a keyword, the set of ads related to the keyword are getting involved in the mechanism. Specifically, the platform collects corresponding bids, decides the number of slots $k$, predicts CTR, and the auction runs.
Motivating example
@@
How would the LLMA be fundamentally different from the SA? We start with illustrative scenarios where a user asks a technical question about computers (Figure 1). Without advertisement, LLM would typically generate a response to address the user’s query. To incorporate advertisements in the generated response and induce user attention, there is a spectrum of possibilities for including ad content: (a) putting the ads outside the response but in the user interface, (b) incorporate the ads within the generated output directly. (a) can be deemed as display ads, and would be relatively easy to handle given the vast amount of techniques in display ads. However, (b) is more similar to search ad or native ad. [^3] Typically, search and native ads capture user attention much better than display ads, e.g., viewed $\sim 50\%$ more than display ads (Outbrain, 2023). We will focus on the approach (b), which will entail fundamental challenges that have not arisen in the traditional SA.  [^4] Also, in case the ads are included within the generated output, there are several possibilities such that the ads may simply replace one of the element in case they are written as a list of elements (e.g., Figure 5, denoted by structured output), or the ads itself is incorporated within the original text/figure without hurting/changing the context of the original output (e.g., Figure 1, denoted by unstructured output). We focus on the case of unstructured output, as this would be broadly applicable in various scenarios, and the structured output may be able to handled with the standard SA framework.
SA versus LLMA
Recall the process of bidding, output generation, prediction, and auction in the SA mentioned before, and imagine implementing those modules for LLMA.
For the bidding module, how could the advertisers write bid? What would they actually bid for? This is not straightforward to answer, since the user query in LLM cannot be simply specified as a keyword. Further, given that the marketing impact will be significantly dependent to how the LLM incorporate the ad in the output, it is not even clear what is the advertiser’s value for being included in the ad. In fact, the advertiser’s value would be dependent to the generated output, then again, how could the advertiser reflect their willingness-to-pay with respect to the output, which might not be accessible in advance? Even further, how can we generate output that smoothly incorporates the ad without hurting the user experience while satisfying the advertiser?
For prediction module, most SA run an online learning algorithm to update the ad’s feature vector with respect to the user context (McMahan et al., 2013). This was possible because the ad images, hyperlinks, and more generally how they appear in the SERP remain the same across many user interactions. LLM, however, would incorporate ad in a very different manner for each query, e.g., see Figure 4, then how can the LLMA learn the CTR fundamentally? Also, since the ads are merged into the generated output, it significantly affects the user experience, then how can we guarantee the user satisfaction and measure them?
Finally, which kind of auction format should LLMA run? How can LLMA adapt for advertising multiple ads in a single output?
All these questions are not straightforward to answer, yet, have not been formally discussed in any literature to the extent of our knowledge.
Our contribution
@@
In this paper, we introduce a vast amount of research questions to operate LLMA in practice. We first discuss several desiderata to run LLMA and propose a seemingly plausible framework to meet the requirements. Similar to SA, our framework consists of four modules, though the implementation of each module will be very different from SA: (i) modification in which the original output of LLM is modified; (ii) bidding that advertisers utilize to bid on the modified outputs; (iii) prediction in which LLMA computes required information about advertisements; and (iv) auction in which the advertises compete and the final output is selected. We discuss different designs and technical challenges of each of those modules, along with further research challenges to implement the system in a practical manner.  [^5] We first focus on presenting a single advertisement in the LLM output, however, generalization of the proposed framework to incorporate multiple ads at once is discussed in Section 5.4. We correspondingly raise several research questions on different design choices for each module. Although we believe that our framework may be grounded as an initial stepping stone for future research, our primary prupose is to discuss research questions that should be addressed for the practical operation of LLM advertisement.
Finally, as an independent interest, in Section 6, we discuss how LLM (more generally generative model) can empower the online advertisement system to tailor the content of ads itself to individual users, leveraging their specific contexts to generate more appealing content, which is called as dynamic creative optimization (DCO) or responsive ad in the literature. The incorporation of DCO into the SA, especially for LLMA, would bring further challenges in each module and we propose corresponding research questions therein.
Figure 2: An example of ChatGPT 4 providing different output tailored to the user context. We keep the context provided in Figure 1, but asks again to advertise again to the specified user segment.
@@ 2 Related works
Here we present related works on LLM, online advertisement system, and their intersection. Further related works can be found in Appendix A.
@@
Advancements in AI, NLP, and conversational agents, driven by Transformer architecture (Vaswani et al., 2017), have given rise to models like GPT-3 (Brown et al., 2020) and BERT (Devlin et al., 2019). These models revolutionize chatbots, enabling context-aware, human-like interactions across diverse domains (Abd-Alrazaq et al., 2020; Nicolescu and Tudorache, 2022). Everyday use of language models has led researchers to investigate the content generated by these models to ensure that they do not hallucinate (Guerreiro et al., 2023; Ji et al., 2023; Li et al., 2023; Zhang et al., 2023) in their outputs, and do not generate harmful or biased content (Liang et al., 2021; Navigli et al., 2023; Kirk et al., 2021; Shen et al., 2023; Weidinger et al., 2021; Liu et al., 2023b). In fact, trustworthy of LLMs is actively studying by researchers (Liu et al., 2024).
@@ 3 Requirements
In this section, we outline our expectations for LLMA, which should be evaluated based on the following criteria:
3.1 System requirement
@@
LLMA needs to associate with advertisers, making the entire system vulnerable to harmful behaviors from them. To mitigate that, LLMA should consider every possible aspect of the behavior of the advertisers, ensuring the reliability and alignment of the system (Hendrycks et al., 2020).
@@ 3.2 Player incentive
When an advertisement is added to LLM output, it’s crucial to maintain a high level of quality in the generated content. Users do not wish to encounter an excessive amount of advertising in the generated responses, especially when it doesn’t align with their search or interests. Advertisement involvement should not significantly degrade LLM output, as doing so would lead to reduced user satisfaction, thereby decreasing the user retention. For instance, in modern online advertisement system, this is often handled using floor prices to filter out irrelevant ads to maintain the quality of the user experience. Similarly for LLM service, we need to ensure the final modified output including advertisements can still be a high-quality response to the user, not deviating significantly from what LLM originally would generate.
@@
As advertisers pay LLMA to include advertisements in their outputs, they expect appropriate exposure of ads in them. Whether their products, services, or any other form of advertising, they expect these items to be showcased in a compelling and engaging manner within the output. This presentation should be interesting for users, ultimately driving revenue for the advertisers.
Platform revenue
@@
Obtaining revenue is the primary goal of operating LLMA, thus, LLMA should ensure that the total additional cost of advertisement is compensated by the revenue obtained from advertisers. It is worth noting that advertisements may potentially reduce the overall number of users engaging with the system, which could have adverse effects on LLM itself.
4 Framework for LLMA
Figure 3: Overall framework of LLMA.
@@ 4 Framework for LLMA
In this section, we present overall framework for LLMA. We introduce several feasible types of architectures for LLMA, each of which has its own advantages and disadvantages that will be discussed together. Mainly, we focus on a scenario in which a user provides a query $q$ to the LLM, and the original output by LLM is given by $X$. Further, a context $c$ captures a variety of features that are relevant to the advertisement recommendation, e.g., history of the previous queries, user segment, region, and date. Although the number of advertisers (bidders) varies from time to time, at the moment the user inputs the query $q$, we suppose that there are $n$ advertisers indexed by $V=\{\text{adv}_{1},\text{adv}_{2},\ldots,\text{adv}_{n}\}$. For simplicity, we assume that each advertiser $\text{adv}_{i}$ is equipped with a single advertisement (ad) he/she wants to post. We denote $\text{adv}_{i}$’s ad by $\text{ad}_{i}$ for $i\in[n]$, and let $D=\{\text{ad}_{1},\ldots,\text{ad}_{n}\}$.
@@
Overall, we divide the LLMA as follows based on their functionalities [^8] These functionalities may be distributed to multiple market players such as supply side platform, display side platform, or ad exchange, as does in the current online ad eco-system. into $4$ modules: (i) output modification, (ii) bidding, (iii) prediction, and (iv) auction. This overall framework is illustrated in Figure 3. In short,
  1. 1.
    The user asks a query $q$ to the system along with a context $c$;
  2. 2.
    The LLMA generates the original output $X$;
  3. 3.
    The modification module generates specific output $X_{i}$ for each of the advertisers;
  4. 4.
    The bidding module generates corresponding bid $b_{i}$. At the same time, the prediction module predicts the quality of the output $X_{i}$ by measuring the user satisfaction rate $\text{sr}_{i}$ and click-through-rate $\text{ctr}_{i}$;
  5. 5.
    Finally, the auction module runs auction to determine the final output and corresponding payment amount to charge the advertiser.
For the rest of the section, we will explain each module and its responsibility/functionality in a sequential manner.
4.1 Modification module
@@
The modification module takes responsibility for generating modified output based on corresponding ads. This module takes the pair of $(q,X,c)$ and a set of advertisements $D$ as the input, and returns $(X_{i})_{i\in[n]}$ where $X_{i}$ denotes the modified output for $\text{ad}_{i}$.
@@
Overall, we consider two approaches based on who generates the modified outputs. 999 For either cases, one can generate the modified output by giving an additional query to the LLM as presented in Figure 1.
@@
    In the LLMA modification model, LLMA directly generates the modified output.
Comparison to SA
Note that the standard SA does not have modification module explicitly as it is trivial to incorporate ads in each slot. Thus, we emphasize that is the challenge that uniquely appears in LLMA.
4.2 Bidding module
The bidding module generates bid based on the modified outputs. The input to the bidding module is the query $q$, context $c$, and modified outputs $(X_{i})_{i\in[n]}$, and the output will be a bid $(b_{i})_{i\in[n]}$, where $b_{i}$ is the bid for $ad_{i}$. Similar to the generation model, we consider two models for the bidding module.
@@ 4.2 Bidding module
The static bidding model would be able to be operated by extracting some keywords from the user query and run the bidding based on the keywords while asking advertisers to bid on each keyword, similar to the SA. In Section 5, we discuss several extensions of these bidding models, thereby allowing more flexibility for advertisers to customize their advertisement strategy by bidding.
@@
Dynamic bidding model may be of interest to the advertisers who have their own measures to estimate the quality of the modified output and thus want to customize their bid in a way that might not be captured by the static bidding. For example, in extreme, suppose that LLMA may only consider click-through-rate of the modified output and does not care about the user satisfaction rate or how much the inserted advertisement relates to the context of the output. In this case, the advertiser may want to customize their bid by measuring the quality of the output by itself, since he may be afraid of the system inserting its ads in an awkward manner.
Comparison to SA
Note that in SA, since the context of user query in search engine is more explicit, and can typically abstracted as several keywords, advertisers can safely bid on each keyword. In LLMA, however, it might be difficult to extract proper keywords from the query as the query itself tends to be much longer than the traditional SA due to the flexibility of LLM. Further, the generated output significantly affects the marketing impact of the ads in LLMA, whereas it is typically assumed (and validated) to be independent from other ads / contents shown within SERP in the standard SA. Finally, the dynamic bidding model further exhibits unique challenges of requiring advertisers to dynamically adjust the bid with respect to the output.
4.3 Prediction Module